For the subspace W= the plane 3x-y+4z-0, of R
1
Expert's answer
2021-07-22T04:50:52-0400

The normal to the plane has the direction "(3, -1, 4)."

Therefore, the basis must consist of two linearly independent vectors that are orthogonal to it.


"3a_1-b_1+4c_1=0"

"3a_2-b_2+4c_2=0"

We consider the set "\{(0, 4, 1), (1, 3, 0)\}."

We can see that for any scalar "k,"


"(0, 4, 1)\not=k (1, 3, 0)"

Therefore, they are linearly independent.

Also,


"(0, 4, 1)\cdot(3, -1, 4)=0"

"(1, 3, 0)\cdot(3, -1, 4)=0"

Therefore, the basis is "\{(0, 4, 1), (1, 3, 0)\}."